what is the temperature coefficient of resistance of a piece of wire having a resistance of 12. 85 ohm at 30 degrees celsius and 6.85 ohm at 0 degrees celcius. what will be its resistance at 50 degrees celsius.


1
Expert's answer
2021-05-04T12:08:38-0400

"\rho=\rho_0(1+\alpha t)"


"12.85=\rho_0(1+\alpha \cdot30)"


"6.85=\rho_0(1+\alpha \cdot0)\to\rho_0=6.85\ (\Omega)" . So, we have


"\alpha=(12.85/6.85-1)/30=0.0292\ (1/°C)" . Answer


"\rho_{50}=6.85(1+0.0292\cdot50)=16.85\ (\Omega)" . Answer